OpenStudy (anonymous):

Will medal. If a sample is taken from a material is initially at 25degrees c, what is its final temperature if 150g of the material absorbs 1000j of heat, and the specific heat of the material is .96 j/g degrees c?

OpenStudy (anonymous):

Don't mind the top part. See that's how I did it but I'm got an answer that doesn't match it.....

OpenStudy (anonymous):

oh the picture only shows up on the app?

OpenStudy (anonymous):

@tom982 @mathmale @pooja195 @zelda101

OpenStudy (anonymous):

Do you know what formula we can use here? Read up on heat capacity equations. You've been given heat, mass, heat capacity and starting temperature.

OpenStudy (anonymous):

ik what the formula is, and i worked it out, but when i got the initial time it was -18, and thats not correct

OpenStudy (anonymous):

1000j= 150g x .96 j/g degrees c x ( )

OpenStudy (anonymous):

i multiplied 150 x .96 and got 144 and divided 1000 by 144 and got -18

OpenStudy (anonymous):

so idk where i messed up

OpenStudy (anonymous):

Our equation is \(Q=cm (T_{final}-T_{initial})\) So we have \(1000=0.96*150(x-25)\) Solving for \(x\) gives \(31.9444444444\)

OpenStudy (anonymous):

how tho

OpenStudy (anonymous):

sorry i have to explain my work

OpenStudy (anonymous):

Just a bit of rearranging. \[1000=0.96*150(x-25)\]\[\frac{1000}{0.96*150}=x-25\]\[6.94444444444=x-25\]\[x=31.94444444444\]

OpenStudy (anonymous):

ok thanks, that makes a lot more sense

OpenStudy (anonymous):

No problem, happy to help.